OpenStudy (anonymous):

Will Medal! If 2tanx/ (1-tan^2x) = 1, then x can equal ____. Check all that apply. A) x=5pi/8 + npi B) x=pi/8 + npi C) x=3pi/8 + npi D) x=7pi/8 + npi

OpenStudy (anonymous):

I think it is A and B. But this is me taking a guess. I am terrible at trig. Sorry

OpenStudy (welshfella):

well first cross multiply to give 2 tanx = 1 - tan^2 x now rearrange this quadratic: tan^2 x + 2tanx - 1 = 0 can you factor this?

OpenStudy (anonymous):

Yes. I am thinking you factor to get tan^2x = -2tan(x)+1?

OpenStudy (welshfella):

no factor as you would any quadratic like x^2 + 3x + 2 = (x + 1)(x + 2)

OpenStudy (welshfella):

oh sorry - on second look this will not factor so you need to use the quadratic formula

OpenStudy (anonymous):

So would a = tan^2x b =2tanx and c=-1?

OpenStudy (welshfella):

no a , b and c are the coefficients of tan^2x and tan x so a = 1 , b = 2 and c = -1

OpenStudy (anonymous):

So x = +/- sqrt (2) in the end right?

OpenStudy (anonymous):

*-1 +/- sqrt(2)

OpenStudy (welshfella):

-2 +/- sqrt(4 - 4*1*-1) / 2 = -2 +/- sqrt8 / 2 = (-2 +/- 2 * sqrt2 )/ 2 = -1 +/- sqrt2 yep

OpenStudy (anonymous):

Awesome! So now do I just plug in this to my original equation?

OpenStudy (welshfella):

well they have the answer options in terms of pi so convert everything to decimals so for example -1 + sqrt2 = 0.4142 so find the angle in radians which is the tangent of 0.4142 and see if one of the options gives the the same angle

OpenStudy (welshfella):

eg angle with tangent 0.4142 = 0.39269 radians now check if pi/8 gives the same value

OpenStudy (anonymous):

Yes it does!

OpenStudy (anonymous):

So I'm assuming that is one

OpenStudy (welshfella):

ok so thats one result pi/8 + npi

OpenStudy (welshfella):

now try the other value for tan x

OpenStudy (anonymous):

-1 - sqrt 2 = -2.4142 degrees right?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Ok so then the radians of -2.4142 is -0.042135

OpenStudy (anonymous):

I don't think I did it right

OpenStudy (welshfella):

no the -2.4142 is the value of tan x so you need tan-1(-2.4242)

OpenStudy (anonymous):

Oh ok that would make it -1.1781 which is basically the same as 3pi/8 right?

OpenStudy (anonymous):

So the answers would be pi/8 + npi and 3pi/8 + npi

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Great!! Thank you for taking the time to help me! I feel I really understand now! Thank you!

OpenStudy (anonymous):

Weird... I hit the sumit button and it told me it was 5pi/8 instead of 3/pi/8

OpenStudy (welshfella):

lets check that out

myininaya (myininaya):

\[\frac{2tanx}{1-\tan^2x} = 1 \\ \tan(2x)=1\]

myininaya (myininaya):

\[\tan(u)=1 \text{ when } u=?\]

OpenStudy (welshfella):

ahh i forgot that identity !

OpenStudy (welshfella):

u = pi/4 of course and x will be pi/8

OpenStudy (welshfella):

much easier that way lol!

myininaya (myininaya):

and there is also the u=5pi/4 so yep x=5pi/8

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

-1.1781 is an incorrect result then? - an extraneous result right?

myininaya (myininaya):

does give tan(2x) approx 1 what number are you approximating ?

myininaya (myininaya):

oh -3pi/8 is what you are approximating

OpenStudy (welshfella):

sorry I dont follow you -yes

myininaya (myininaya):

that gives 1

myininaya (myininaya):

when you plug in

OpenStudy (welshfella):

and -3pi/8 is equivalent to 5pi/8

myininaya (myininaya):

is that not included in the solution: \[\frac{5\pi}{8}+n \pi \\ \text{ if } n=-1 \text{ then } \frac{5\pi}{8}-\pi=\frac{-3\pi}{8}\] so I'm not sure why you say it is extraneous

OpenStudy (welshfella):

no - i was wrong its not extraneous

OpenStudy (welshfella):

still i think lehmad learned something along the line...

OpenStudy (anonymous):

Yes I learned a lot! Thank you guys for the help.